# Unveiling the jet angular broadening with photon-tagged jets in high-energy nuclear collisions

### v2

@@ III Results and discussions
Recently, the medium modifications of $\gamma$+jet girth distribution in PbPb collisions $\sqrt{s_{NN}}=5.02$ TeV are measured by the CMS collaboration CMS:2024zjn . The jet girth belongs to a generalized class of observables known as jet angularities Larkoski:2014pca ; KunnawalkamElayavalli:2017hxo , defined as,
$g=\frac{1}{p_{T}^{\rm jet}}\sum_{i\in\rm jet}p_{T}^{i}\Delta R_{i,\rm jet}.$ (8)
@@
where the index $i$ sums over all jet constituents, and $\Delta R_{i,\rm jet}$ represents the angular distance between each particle and the jet axis in the $\eta-\phi$ plane. The girth value quantifies the $p_{T}$ distribution of particles within a jet, weighted by their angular distance from the jet axis. This observable is expected to be sensitive to modifications in the jet angular structure that occur during heavy-ion collisions ALICE:2018dxf ; CMS:2024zjn ; Yan:2020zrz ; Li:2024pfi . The jet-medium interactions, such as the parton scattering, the medium-induced gluon radiation and the medium response, may increase the values of jet girth in heavy-ion collisions compared to those in pp.
@@
Figure 1: (Color online) Normalized girth distributions of $\gamma$+jets in pp $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV calculated with $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (middle panel) as compared to the recent CMS data CMS:2024zjn . The ratios (PbPb/pp) of girth distribution are also shown in the lower panel.
@@
As shown in Fig. 1, we present the theoretical results of the $\gamma$+jet girth distribution for $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (middle panel) in both pp and $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, compared with the CMS data. The lower panels show the ratio PbPb/pp distributions. All selected jets are constructed with the anti-$k_{T}$ algorithm and are required to have a transverse momentum of $p_{T}^{\rm jet}>40$ GeV, with pseudorapidity $|\eta_{\rm jet}|<2$ and jet radius $R=0.2$, while the photon must have $p_{T}^{\gamma}>100$ GeV and $|\eta_{\gamma}|<1.44$. In addition, the selected photon must meet the isolation requirement, which states that the sum of the transverse momentum of all particles within a distance $R=0.4$ around the photon must be less than 5 GeV. Furthermore, the jet and photon should be nearly “back-to-back”, requiring $\Delta\phi_{\gamma,\rm jet}>2\pi/3$. Our theoretical calculations provide a satisfactory description of the girth distribution of the CMS data for both pp and PbPb collisions for the cuts $x_{j\gamma}>0.4$ and $x_{j\gamma}>0.8$ CMS:2024zjn . Notably, a significant difference arises in the ratios of girth distributions between PbPb and pp collisions (PbPb/pp) for the two different $x_{j\gamma}$ cuts. For $x_{j\gamma}>0.4$, the modification of girth is modest, with an enhancement observed between $0.10.08<g<$ 0.1. Since girth quantifies the angular-weighted transverse momentum distribution of jets, an increase in PbPb/pp at a larger girth indicates that jets are broader in PbPb collisions compared to pp collisions. In contrast, for $x_{j\gamma}>0.8$, we observe an enhancement at $g>0.02$ and a clear suppression at $g>0.04$, suggesting that jets are narrower compared to the case of $x_{j\gamma}>0.4$. It prompts the question: how does the $x_{j\gamma}$ selection cut affect the patterns of jet angular structure modifications in nucleus-nucleus collisions? One possible explanation is that the lower $x_{j\gamma}$ cut allows for the inclusion of more significantly quenched jets traversing the QGP, which reduces the selection bias effect CMS:2024zjn . To explore this hypothesis, we conduct a Jet-by-Jet matching procedure to investigate the relationship between selection bias and kinematic cuts in heavy-ion collisions.
@@
In experimental measurements within the field of high-energy heavy-ion collisions, the conventional approach to study nuclear modification effects on jet substructure involves applying identical dynamical constraints during jet reconstruction and selection for both proton-proton and nucleus-nucleus collisions. By comparing the selected jet samples from these two collision systems, one can calculate the nuclear modification effects for the specific jet observables, such as jet girth. However, since the jets in these two samples do not correspond to the same individual jet before and after quenching, the derived nuclear modification effects represent an overall statistical average across a large ensemble of jets. Therefore, it is natural to inquire about how jet-medium interactions affect the substructure of individual jets before and after quenching in nucleus-nucleus collisions. While experimentally tracking the evolution of a single jet through different stages remains highly challenging, such processes can indeed be modeled and analyzed via a jet-by-jet matching in the Monte Carlo simulations as performed in the studies of Refs. Brewer:2021hmh ; Kang:2023ycg . The details are illustrated as follows:
@@
Jet-by-Jet matching: The events selected with suitable experimental kinematic cuts, such as $p_{T}>$ 40 GeV and $|\eta_{jet}|<$ 2.0, in pp collisions are used as the input of jet evolution in PbPb. We can reconstruct the jets in pp and PbPb collisions by the event particles before and after the in-medium evolutions. By calculating the angular distance $\Delta R<R$ between the axis of each jet pair in pp and PbPb, the nearest pair of jets in the $\eta-\phi$ plane are regarded as the matched one before and after quenching. To consider the jets, with $p_{T}>$ 40 GeV initially, dropping down to the cut due to the jet energy loss, we use a lower cut $p_{T}>$ 10 GeV to select the possible candidates in PbPb collisions. This Jet-by-Jet (JBJ) matching method effectively eliminates selection bias, enabling precise tracking of the medium evolution of individual jets in the Monte Carlo simulations. It offers a novel perspective to study jet-medium interactions in high-energy heavy-ion collisions and provides critical insights for interpreting recent experimental measurements of jet substructure modifications.
@@
As shown in Fig. 2, with the help of the Jet-by-Jet matching, we compare the $x_{j\gamma}$ distribution of $\gamma$+jets before and after quenching in $0-30\%$ PbPb collisions when using different $x_{j\gamma}$ cuts, $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (lower panel). The solid line represents the initially selected (unquenched) jets above $x_{j\gamma}$ cut, while the dashed line denotes the corresponding jet sample after the in-medium evolution (quenched). Due to energy loss in the QGP, the $x_{j\gamma}$ distributions of quenched jets shift toward lower $x_{j\gamma}$ values compared to unquenched one. The shaded region represents the jets which have suffered sufficient quenching effect in nucleus-nucleus collisions and have dropped down to the $x_{j\gamma}$ cut. The contribution of the shaded region was usually discarded in the realistic experimental measurements in PbPb collisions, even though they initially have $p_{T}>40$ GeV. Therefore, the fraction of the shaded region quantitatively reflect the influence of selection bias, namely the amount of jets with sufficient quenching effect but are rejected by the selection requirement in the measurement in PbPb collisions. We find that the fraction of the shaded region is $23.9\%$ for $x_{j\gamma}>0.4$ and $58.4\%$ for $x_{j\gamma}>0.8$. This means that when applying the criterion $x_{j\gamma}>0.8$, less than half of the jets survive the event selection in PbPb collisions, thereby excluding a significant number of quenched jets. In contrast, using $x_{j\gamma}>0.4$ includes more jets that experience substantial quenching in PbPb collisions, ultimately resulting in a broader modification of the $\gamma$+jet girth compared to pp, as observed in the CMS measurement CMS:2024zjn .
@@
Figure 3: (Color online) Medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV as compared to the CMS data CMS:2024zjn . In the upper panel, the results without considering the medium-induced gluon radiation and medium response are also presented. In the lower panel, we also compare the girth modification of inclusive jet and $\gamma$+jet with the same $p_{T}$ cut, as well as the case using the Jet-by-Jet matching (JBJ).
@@
Furthermore, measurements of inclusive jets in PbPb collisions indicate a narrowing effect ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , while observations from $\gamma$+jets suggest hints of broadening CMS:2024zjn . It will be of great significance to theoretically explore the differing patterns of substructure modification between these two jet samples within the same collision system. In Fig. 3, we present the girth modification of $\gamma$+jets and inclusive jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. Aside from the tagging requirements for $\gamma$+jets—specifically, $p_{T}^{\gamma}>100$ GeV, $x_{j\gamma}>0.4$, and $|\eta_{\gamma}|<1.44$—all selected jets must have $p_{T}>40$ GeV and $|\eta_{\text{jet}}|<2$. In the upper panel, we first discuss the influence of medium-induced gluon radiation and the medium response on $\gamma$+jets. Compared to the results without gluon radiation, our findings indicate that medium-induced gluon radiation play a significant role in the jet angular broadening in nucleus-nucleus collisions. Additionally, we also find that the medium response slightly enhances the modification in the region of larger girth. In the lower panel, we compare the girth modifications of $\gamma$+jets and inclusive jets. We observe that inclusive jets show suppression at $g>0.05$, which indicates a narrowing modification consistent with previous ALICE measurements in PbPb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2018dxf . This behavior differs from that of $\gamma$+jets. To address this discrepancy, we calculate the girth modification of the jet sample using the Jet-by-Jet matching method, which represents the jet modification in the hot QCD matter without selection bias. Both $\gamma$+jets and inclusive jets show consistent and noticeable broadening at larger girth compared to the initial jets. In other words, jets naturally become broader due to jet-medium interactions in PbPb collisions for both inclusive jets and $\gamma$+jets. We will demonstrate that selection bias plays different roles in $\gamma$+jets and inclusive jets, ultimately leading to the distinct modification patterns observed in experimental measurements.
@@
Figure 4: (Color online) Normalized $p_{T}$ distribution of unquenched and quenched jets for $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV by using the Jet-by-Jet matching procedure.
@@
In Fig. 4, we also compare the $p_{T}$ spectra of the jets before and after quenching in PbPb collisions for both $\gamma$+jets and inclusive jets, obtained using the JBJ matching. By applying the same jet $p_{T}$ cut (specifically, $p_{T}>40$ GeV), we observe that the shape of the $\gamma$+jet $p_{T}$ spectrum before quenching significantly differs from that of the inclusive jets. The $\gamma$+jet spectrum increases gradually with $p_{T}$, peaking near the photon $p_{T}$. In contrast, the inclusive jet distribution is mainly concentrated in the range of $[40,80]$ GeV and decreases rapidly with increasing $p_{T}$. Due to in-medium energy loss, we find that the $p_{T}$ distributions of both $\gamma$+jets and inclusive jets in PbPb collisions shift towards a lower energy region compared to their pp counterparts. The shaded region indicates the jets that do not meet the $p_{T}>40$ GeV requirement due to energy loss in PbPb. Specifically, the fraction of the shaded region is $20.3\%$ for $\gamma$+jets and $63.4\%$ for inclusive jets. In other words, approximately $80\%$ of $\gamma$+jets survive the selection criteria in PbPb, whereas less than $40\%$ of inclusive jets do. This indicates that using jets associated with direct photons can significantly reduce selection bias. It can be understood from two perspectives. First, inclusive jets are mostly distributed near the selection cut (40 GeV), while jets associated with photons cover a broader $p_{T}$ range with a peak near the $p_{T}$ of the triggered photon (100 GeV). As a result, the unique $p_{T}$ spectrum of the $\gamma$+jets provides a much lower probability of jets falling below the cut-off after quenching compared to the inclusive jets. Second, the inclusive jets consist of $60\%$ gluon jets at $p_{T}^{\rm jet}>40$ GeV, whereas the $\gamma$+jet process is quark-jet dominated (above $70\%$ jets are quark-initiated jets at $p_{T}^{\rm jet}>40$ GeV for $p_{T}^{\rm\gamma}>100$ GeV). Consequently, the inclusive jet may lose more energy, resulting in a lower survival rate to pass the selection criteria in PbPb.
@@
In addition to the higher survival probability of $\gamma$+jets compared to inclusive jets, we are also interested in understanding the differences in the quenching strength experienced by the surviving samples for these two types of jets within the QGP. Since quenching strength is not a specifically defined physical quantity, we assume that jet energy loss can effectively quantify the intensity of the interaction between the jet and the medium. Therefore, we estimate the event-averaged $p_{T}$ loss, denoted as $\langle\Delta p_{T}\rangle=\langle p_{T}^{\rm init}-p_{T}^{\rm fin}\rangle_{\text{evt}}$, for both $\gamma$+jets and inclusive jets as a function of final jet $p_{T}$ ($p_{T}^{\rm fin}$) in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, using Jet-by-Jet matching as illustrated in Fig. 5. In the left panel, we plot the $p_{T}$ loss of $\gamma$+jets and inclusive jets as a function of their initial transverse momentum. It can be observed that $\langle\Delta p_{T}\rangle$ for both $\gamma$+jets and inclusive jets increases with $p_{T}^{\rm init}$, with the inclusive jets showing larger magnitudes. This aligns with expectations, as inclusive jets contain a larger fraction of gluon jets, while $\gamma$+jets are predominantly comprised of quark jets. In the right panel of Fig. 5, we also plot the $\langle\Delta p_{T}\rangle$ for $\gamma$+jets and inclusive jets as a function of $p_{T}^{\rm fin}$. Surprisingly, we find that $\langle\Delta p_{T}\rangle$ for $\gamma$+jets is significantly larger than that for inclusive jets in the range of $8040<p_{T}^{\rm fin}<$ 80 GeV. This implies that selected $\gamma$+jets with $p_{T}>40$ GeV in PbPb collisions have statistically experienced stronger quenching than the inclusive jets. Compared to inclusive jets, the specific initial $p_{T}$ distribution of $\gamma$+jets, which includes a substantial number of jets with much higher $p_{T}$ than the selection threshold, allows for the possibility that jets experiencing significant quenching can still survive in the selection in PbPb collisions. Therefore, we have quantitatively demonstrated that $\gamma$+jets provide unique and significant advantages in reducing selection bias and effectively collecting sufficiently quenched jets in PbPb collisions compared to inclusive jets.
@@
The discussions above highlight the connection between selection bias and the choice of the threshold $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. To offer theoretical guidance for searching more pronounced angular broadening effects of $\gamma$+jets in experimental settings, we have conducted calculations that explore the sensitivities of jet girth modifications to the $x_{j\gamma}$ and $p_{T}^{\gamma}$ cuts, as illustrated in Fig. 6. In the upper panel, we calculate the girth modification of $\gamma$+jet in PbPb collisions with varying $x_{j\gamma}$ cuts. It is evident that when $x_{j\gamma}>0.2$, there is more noticeable angular broadening because more quenched jets are accepted in the selection process in PbPb. As the $x_{j\gamma}$ cut increases from 0.2 to 0.8, the increasing influence of selection bias leads to a transition from enhancement at $g>0.06$ to suppression, namely the modification pattern of jet changes from broadening to narrowing, consistent with the trend observed in the CMS data CMS:2024zjn . Note that we have applied a lower $p_{T}$ cut of 20 GeV to reconstruct the jets, ensuring that the variation of $x_{j\gamma}$ down to 0.2 remains accessible. Additionally, the lower panel presents calculations of $\gamma$+jet girth modification in PbPb collisions with different $p_{T}^{\gamma}$ cuts. We find that applying a higher $p_{T}$ cut for the photon can also help mitigate the influence of selection bias and result in more pronounced girth modifications of $\gamma$+jets in PbPb collisions.
Figure 7: (Color online) Medium modification of the groomed radius distribution of $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of $\gamma$+jet by the CMS CMS:2024zjn . We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
To provide a comprehensive analysis, we will also examine other jet substructure observables and highlight their different modification patterns for $\gamma$+jets and inclusive jets, similar to our findings regarding jet girth. In Fig. 7, we compare the medium modification of the groomed radius ($R_{g}$) distribution for $\gamma$+jets and inclusive jets in $0-30\%$ Pb–Pb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp refers to the ratio of normalized $R_{g}$ distributions in Pb–Pb to pp collisions. Here, $R_{g}$ represents the opening angle between the two hardest subjets (denoted as $1$ and $2$) that satisfy the Soft Drop condition Larkoski:2014wba ,

### v4

@@ III Results and discussions
Recently, the medium modifications of $\gamma$+jet girth distribution in PbPb collisions $\sqrt{s_{NN}}=5.02$ TeV are measured by the CMS collaboration Hayrapetyan et al. (2025b). The jet girth belongs to a generalized class of observables known as jet angularities Larkoski et al. (2014a); Kunnawalkam Elayavalli and Zapp (2017), defined as,
$g=\frac{1}{p_{T}^{\rm jet}}\sum_{i\in\rm jet}p_{T}^{i}\Delta R_{i,\rm jet},$ (7)
@@
where the index $i$ sums over all jet constituents, and the $\Delta R_{i,\rm jet}$ represents the angular distance between each particle and the jet axis in the $\eta-\phi$ plane. The jet axis is determined by summing the momenta of all constituent particles of the anti-$k_{T}$ jet using the “E-scheme” clustering algorithm Cacciari et al. (2008). The girth value quantifies the $p_{T}$ distribution of particles within a jet, weighted by their angular distance from the jet axis. This observable is expected to be sensitive to modifications in the jet angular structure that occur during heavy-ion collisions Acharya et al. (2018); Hayrapetyan et al. (2025b); Yan et al. (2021); Li et al. (2025). The jet-medium interactions, such as the parton scattering, medium-induced gluon radiation, and medium response, may increase the values of jet girth in heavy-ion collisions compared to those in pp.
@@
Figure 1: (Color online) Normalized girth distributions of $\gamma$+jets in pp $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV calculated with $x_{j\gamma}>$ 0.4 (upper panel) and $x_{j\gamma}>$ 0.8 (middle panel) as compared to the recent CMS data Hayrapetyan et al. (2025b). The ratios (PbPb/$p$) of girth distribution are also shown in the lower panel.
@@
As shown in Fig. 1, we present the theoretical results of the $\gamma$+jet girth distribution for $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (middle panel) in both pp and $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, compared with CMS data. The lower panels show the ratio PbPb/pp distributions. All selected jets are constructed with the anti-$k_{T}$ algorithm and are required to have a transverse momentum of $p_{T}^{\rm jet}>40$ GeV, with pseudorapidity $|\eta_{\rm jet}|<2$ and jet radius $R=0.2$, while the photon must have $p_{T}^{\gamma}>100$ GeV and $|\eta_{\gamma}|<1.44$. In addition, the selected photon must meet the isolation requirement, which states that the sum of the transverse momentum of all particles within a distance $R=0.4$ around the photon must be less than 5 GeV. Furthermore, the jet and photon should be nearly “back-to-back”, requiring $\Delta\phi_{\gamma,\rm jet}>2\pi/3$. The medium evolution was simulated over a statistically significant ensemble of $10^{7}$ jet events, thereby guaranteeing that the relative statistical error in our theoretical calculations is negligible, quantified at less than $1\%$. This high precision provides a solid foundation for the subsequent analysis and discussion. Our theoretical calculations provide a satisfactory description of the CMS girth distribution data in both pp and PbPb collisions for the cuts $x_{j\gamma}>0.4$ and $x_{j\gamma}>0.8$ Hayrapetyan et al. (2025b). Notably, the ratio of girth distributions between PbPb and pp collisions (PbPb/pp) differs significantly for the two $x_{j\gamma}$ cuts. For $x_{j\gamma}>0.4$, the modification of girth is modest, with an enhancement observed between $0.08<g<0.1$. Since girth quantifies the angular-weighted transverse momentum distribution of jets, an increase in PbPb/pp at a larger girth indicates that jets are broader in PbPb collisions compared to pp collisions. In contrast, for $x_{j\gamma}>0.8$, we observe an enhancement at $g>0.02$ and a clear suppression at $g>0.04$, suggesting that jets are narrower compared to the condition $x_{j\gamma}>0.4$.
[…]
@@
In experimental measurements within the field of high-energy heavy-ion collisions, the conventional approach to studying the nuclear modification effects on jet substructure involves applying identical dynamical constraints during jet reconstruction and selection for both proton-proton and nucleus-nucleus collisions. By comparing the selected jet samples from these two collision systems, one can calculate the nuclear modification effects for the specific jet observables, such as jet girth. However, since the jets in these two samples do not correspond to the same individual jet before and after quenching, the derived nuclear modification effects represent an overall statistical average across a large ensemble of jets. Therefore, it is natural to inquire about how jet-medium interactions affect the substructure of individual jets before and after quenching in nucleus-nucleus collisions. While experimentally tracking the evolution of a single jet through different stages remains highly challenging, such processes can indeed be modeled and analyzed via a jet-by-jet matching in the Monte Carlo simulations as performed in the studies of Refs. Brewer et al. (2022); Kang et al. (2025a). The details are illustrated as follows:
@@
Jet-by-Jet matching: The events selected with suitable experimental kinematic cuts, such as $p_{T}>$ 40 GeV and $|\eta_{jet}|<$ 2.0, in pp collisions are used as the input of jet evolution in PbPb. We can reconstruct the jets in pp and PbPb collisions by the event particles before and after the in-medium evolutions. We calculate the angular distance $\Delta R<R$ between the axes of each jet pair in pp and PbPb events. The jet pair with the smallest $\Delta R$ is then regarded as the matched one before and after quenching. To consider the jets, with $p_{T}>$ 40 GeV initially, dropping down to the cut due to jet energy loss, we use a lower cut $p_{T}>$ 10 GeV to select the possible candidates in PbPb collisions. This Jet-by-Jet (JBJ) matching method effectively eliminates selection bias, enabling precise tracking of the medium evolution of individual jets in the Monte Carlo simulations. It offers a novel perspective to study jet-medium interactions in high-energy heavy-ion collisions and provides critical insights for interpreting recent experimental measurements of jet substructure modifications.
@@
As shown in Fig. 2, with the help of the Jet-by-Jet matching, we compare the $x_{j\gamma}$ distribution of $\gamma$+jets before and after quenching in $0-30\%$ PbPb collisions when using different $x_{j\gamma}$ cuts, $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (lower panel). The solid line represents the initially selected (unquenched) jets above $x_{j\gamma}$ cut, while the dashed line denotes the corresponding jet sample after the in-medium evolution (quenched). Due to energy loss in QGP, the $x_{j\gamma}$ distributions of quenched jets shift toward lower $x_{j\gamma}$ values compared to the unquenched ones. The shaded region represents the jets that have suffered sufficient quenching effect in nucleus-nucleus collisions and have dropped down to the $x_{j\gamma}$ cut. The contribution of the shaded region was usually discarded in the realistic experimental measurements in PbPb collisions, even though they initially have $p_{T}>40$ GeV. Therefore, the fraction of the shaded region quantitatively reflects the influence of selection bias, namely the amount of jets with sufficient quenching effect but are rejected by the selection requirement in the measurement in PbPb collisions. We find that the fraction of the shaded region is $23.9\%$ for $x_{j\gamma}>0.4$ and $58.4\%$ for $x_{j\gamma}>0.8$. It means that when applying the criterion $x_{j\gamma}>0.8$, less than half of the jets survive the event selection in PbPb collisions, thereby excluding a significant number of quenched jets. In contrast, using $x_{j\gamma}>0.4$ includes more jets that experience substantial quenching in PbPb collisions, ultimately resulting in a broader modification of the $\gamma$+jet girth compared to pp, as observed in the CMS measurement Hayrapetyan et al. (2025b).
@@
Figure 3: (Color online) Medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV as compared to the CMS data Hayrapetyan et al. (2025b). In the upper panel, the results without considering the medium-induced gluon radiation and medium response are also presented. In the lower panel, we also compare the girth modification of inclusive jet and $\gamma$+jet with the same $p_{T}$ cut, as well as the case using the Jet-by-Jet matching (JBJ).
@@
Furthermore, measurements of inclusive jets in PbPb collisions indicate a narrowing effect Acharya et al. (2022); Aad et al. (2023a); Acharya et al. (2018, 2023); Ehlers (2022); Aad et al. (2023b), while observations from $\gamma$+jets suggest hints of broadening Hayrapetyan et al. (2025b). It will be of great significance to explore the differing patterns of substructure modification between these two jet samples within the same collision system. In Fig. 3, we present the girth modification of $\gamma$+jets and inclusive jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. Aside from the tagging requirements for $\gamma$+jets—specifically, $p_{T}^{\gamma}>100$ GeV, $x_{j\gamma}>0.4$, and $|\eta_{\gamma}|<1.44$—all selected jets must have $p_{T}>40$ GeV and $|\eta_{\text{jet}}|<2$. In the upper panel, we first discuss the influence of medium-induced gluon radiation and medium response on $\gamma$+jets. Compared to the results without gluon radiation, our findings indicate that medium-induced gluon radiation plays a significant role in the jet angular broadening in nucleus-nucleus collisions. Additionally, we also find that medium response slightly enhances the modification in the region of larger girth. In the lower panel, we compare the girth modifications of $\gamma$+jets and inclusive jets. We observe that inclusive jets show suppression at $g>0.05$, which indicates a narrowing modification consistent with previous ALICE measurements in PbPb collisions at $\sqrt{s_{NN}}=2.76$ TeV Acharya et al. (2018). This behavior differs from that of $\gamma$+jets. To address this discrepancy, we calculate the girth modification of the jet sample using the Jet-by-Jet matching method, which represents the jet modification in the hot QCD matter without selection bias. Both $\gamma$+jets and inclusive jets show consistent and noticeable broadening at larger girth compared to the initial jets. In other words, jets intrinsically become broader due to jet-medium interactions in PbPb collisions for both inclusive jets and $\gamma$+jets. We will demonstrate that selection bias plays different roles in $\gamma$+jets and inclusive jets, ultimately leading to the distinct modification patterns observed in experimental measurements.
@@
Figure 4: (Color online) Normalized $p_{T}$ distribution of unquenched and quenched jets for $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV by using the Jet-by-Jet matching procedure. The gluon-jet fractions in inclusive jet and $\gamma+$jet samples are also estimated in the lower panel.
@@
In Fig. 4, we also compare the $p_{T}$ spectra of the jets before and after quenching in PbPb collisions for both $\gamma$+jets and inclusive jets, obtained using the JBJ matching. By applying the same jet $p_{T}$ cut (specifically, $p_{T}>40$ GeV), we observe that the shape of the $\gamma$+jet $p_{T}$ spectrum before quenching significantly differs from that of the inclusive jets. The $\gamma$+jet spectrum increases gradually with $p_{T}$, peaking near the photon $p_{T}$. In contrast, the inclusive jet distribution is mainly concentrated in the range of $[40,80]$ GeV and decreases rapidly with increasing $p_{T}$. Due to in-medium energy loss, we find that the $p_{T}$ distributions of both $\gamma$+jets and inclusive jets in PbPb collisions shift towards a lower energy region compared to their pp counterparts. The shaded region indicates the jets that do not meet the $p_{T}>40$ GeV requirement due to energy loss in PbPb. Specifically, the fraction of the shaded region is $20.3\%$ for $\gamma$+jets and $63.4\%$ for inclusive jets. In other words, approximately $80\%$ of $\gamma$+jets survive the selection criteria in PbPb, whereas less than $40\%$ of inclusive jets do. It indicates that using jets associated with direct photons can significantly reduce selection bias, as can be understood from two perspectives. First, inclusive jets are mostly distributed near the selection cut (40 GeV), while jets associated with photons cover a broader $p_{T}$ range with a peak near the $p_{T}$ of the triggered photon (100 GeV). As a result, the unique $p_{T}$ spectrum of the $\gamma$+jets provides a much lower probability of jets falling below the cut-off after quenching compared to the inclusive jets. Second, the inclusive jet sample contains a significant fraction of gluon jets, whereas the $\gamma+$jet process is dominated by quark jets. As shown in the lower panel of Fig. 4, the initial fraction of gluon jets in the inclusive jet sample is approximately $60\%$. This fraction decreases after quenching because the gluon’s larger color charge leads to greater energy loss compared to quarks. In contrast, the gluon jet fraction in the $\gamma+$jet sample is significantly lower (around $10-20\%$ initially) and remains at about $10\%$ after quenching. Consequently, inclusive jets may lose more energy on average, resulting in a lower survival rate for passing the selection criteria in PbPb collisions.
@@
In addition to the higher survival probability of $\gamma$+jets compared to inclusive jets, we are also interested in understanding the differences in the quenching strength experienced by the surviving samples for these two types of jets within QGP. Since quenching strength is not a specifically defined physical quantity, we assume that jet energy loss can effectively quantify the intensity of the interaction between the jet and the medium. Therefore, we estimate the event-averaged $p_{T}$ loss, denoted as $\langle\Delta p_{T}\rangle=\langle p_{T}^{\rm init}-p_{T}^{\rm fin}\rangle_{\text{evt}}$, for both $\gamma$+jets and inclusive jets as a function of final jet $p_{T}$ ($p_{T}^{\rm fin}$) in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, using Jet-by-Jet matching as illustrated in Fig. 5. In the left panel, we plot the $p_{T}$ loss of $\gamma$+jets and inclusive jets as a function of their initial transverse momentum. It can be observed that $\langle\Delta p_{T}\rangle$ for both $\gamma$+jets and inclusive jets increases with $p_{T}^{\rm init}$, with the inclusive jets showing larger magnitudes. It aligns with expectations, as inclusive jets contain a larger fraction of gluon jets, while $\gamma$+jets are predominantly comprised of quark jets. In the right panel of Fig. 5, we also plot the $\langle\Delta p_{T}\rangle$ for $\gamma$+jets and inclusive jets as a function of $p_{T}^{\rm fin}$. Surprisingly, we find that $\langle\Delta p_{T}\rangle$ for $\gamma$+jets is significantly larger than that for inclusive jets in the range of $40<p_{T}^{\rm fin}<80$ GeV. It implies that selected $\gamma$+jets with $p_{T}>40$ GeV in PbPb collisions have statistically experienced stronger quenching than inclusive jets. Compared to inclusive jets, the specific initial $p_{T}$ distribution of $\gamma$+jets, which includes a substantial number of jets with much higher $p_{T}$ than the selection threshold, allows for the possibility that jets experiencing significant quenching can still survive in the selection in PbPb collisions. Therefore, we have quantitatively demonstrated that $\gamma$+jets provide unique and significant advantages in reducing selection bias and effectively collecting sufficiently quenched jets in PbPb collisions compared to inclusive jets.
@@
The discussions above highlight the connection between selection bias and the choice of the threshold $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. To offer theoretical guidance for searching for more pronounced angular broadening effects of $\gamma$+jets in experimental settings, we have conducted calculations that explore the sensitivities of jet girth modifications to the $x_{j\gamma}$ and $p_{T}^{\gamma}$ cuts, as illustrated in Fig. 6. In the upper panel, we calculate the girth modification of $\gamma$+jet in PbPb collisions with varying $x_{j\gamma}$ cuts. It is evident that when $x_{j\gamma}>0.2$, more pronounced angular broadening is observed because more quenched jets are accepted in the selection process in PbPb. As the $x_{j\gamma}$ cut increases from 0.2 to 0.8, the increasing influence of selection bias leads to a transition from enhancement at $g>0.06$ to suppression, namely the modification pattern of the jet changes from broadening to narrowing, consistent with the trend observed in CMS data Hayrapetyan et al. (2025b). Note that we have applied a lower $p_{T}$ cut of 20 GeV to reconstruct the jets, ensuring that the variation of $x_{j\gamma}$ down to 0.2 remains accessible. Additionally, the lower panel presents calculations of $\gamma$+jet girth modification in PbPb collisions with different $p_{T}^{\gamma}$ cuts. We find that applying a higher $p_{T}$ cut for the photon can also help mitigate the influence of selection bias and result in more pronounced girth modifications of $\gamma$+jets in PbPb collisions.
Figure 7: (Color online) Medium modification of the groomed radius distribution of $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of $\gamma$+jet by the CMS Hayrapetyan et al. (2025b). We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
To provide a comprehensive analysis, we will also examine other jet substructure observables and highlight their different modification patterns for $\gamma$+jets and inclusive jets, similar to our findings regarding jet girth. In Fig. 7, we compare the medium modification of the groomed radius ($R_{g}$) distribution for $\gamma$+jets and inclusive jets in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp refers to the ratio of normalized $R_{g}$ distributions in PbPb to pp collisions. Here, $R_{g}$ represents the opening angle between the two hardest subjets (denoted as $1$ and $2$) that satisfy the Soft Drop condition Larkoski et al. (2014b),